OpenStudy (anonymous):

A line passes through (2, –1) and (8, 4).Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

first find slope (m). m=y2−y1x2−x1 equation of line: (y−y1)=m(x−x1) assume: (x1,y1) (2,-1) ........................................ (x2,y2) (8,4)